# Entanglement of edge modes in (very) strongly correlated topological insulators

### 21 Dec 2023 v1

@@ I Introduction
Albeit, topological insulators have been classified [10, 11] even in the presence of interaction, defining an order parameter is still not well understood. In the Su-Schrieffer-Heeger (SSH) model [12, 13, 14], the prototypical example of 1-dimensional topological insulators, two distinct topological phases are attributed due to staggered hopping amplitudes. In the non-trivial topological phase, there are two zero energy edge modes. Motivated by the fact that the existence of edge modes is a prominent signature of a topologically non-trivial phase, the entanglement entropy of edge modes has been studied in SSH or similar models extensively in absence of interaction [15, 16, 17, 18, 19, 20, 21, 22, 23]. Following this cue, we probe a simple model with a non-trivial topological phase structure, i.e. a bi-layer SSH [24, 25, 26, 27, 28, 29] model via studying its entanglement features for a wide range of interaction strengths.
@@
We study the system for the extremely large value of interaction strength more closely in section V. At very large coupling, if one neglects the hopping terms, the theory becomes ultra-local. Ultra-local models like these are called Carrollian field theories, which have recently gained interest in condensed matter [53, 54], fluid dynamics [55, 56, 57, 58], high energy [59, 60, 61] and gravitational physics [62, 63] for diverse inquiries. In particular, fermionic models with these symmetries have been dealt with in detail in [64, 65, 66]. For ultra-local systems like this, the von Neuman entropy is expected to drop to zero, which is also supported by DMRG calculations. However, if one considers including the hopping terms, albeit perturbatively small, their effects still characterize the topological phase of the system, depending upon the values of the parameters, and that is reflected in the ground state entanglement of the system as illustrated in Fig. 9.
@@ III.3 Symmetry resolved entanglement
where $\mathcal{Z}_{n}(q)$ is given by the $n$-sheetd replica copy of the density matrix trace projected on the $q$-charge sector:
@@ IV.1.3 Discussion
has zero energy with respect to the free theory Hamiltonian, in the non-trivial topological phase. The $A$ coefficients appearing in (16) are complex numbers subject to the condition that the state is normalized. Hence, the edge-state is 4-fold degenerate. For the same reason, the ground state,
@@
is also degenerate. Here, $b_{k}$ are all the finite, negative energy modes (holes). The 4-bilinears $\alpha^{\dagger}_{i,\uparrow}\alpha^{\dagger}_{j,\downarrow}$ are bosonic in nature. Hence, the degenerate edge state space is an orbit of $SU(4)$ action on linear combinations of 4-bosons.
At the thermodynamic limit, $\alpha_{i}$ are exactly localized at the edges, and the maximally entangled edge state is the one when all the 4 $A$’s appearing in (16) are equal, modulo $U(1)$ phases. In that case the edge state is a couple of Bell pairs. This gives an entanglement entropy of $2\ln 2$.
@@ V Entanglement signatures at extreme coupling
Choosing a set of functions $f$ as Dirac delta functions supported at each supported at points on the spatial line, we see that we have a conserved operator for each degree of freedom located at each spatial point, which mutually commutes. In the discretized version, these operators are $U\psi^{\dagger}_{\uparrow}\psi_{\uparrow}\psi^{\dagger}_{\downarrow}\psi_{\downarrow}$, ie the Hamiltonian density at each point. Hence, the system is exactly solvable. In fact, in this model, the local $U(1)$ phase transformations:
@@
However, the ground state being so much degenerate, the meaning of bi-partition entanglement becomes ambiguous to a large extent. For example, the vacuum state, which also is a member of the 0 energy subspace, is entanglement-free. This is a prototypical example of an entanglement-free ground state of ultra-local conformal theories [74], at least in cases where the ground state is unique. On the other hand, one can construct a linear combination of ${}^{L}C_{L/2}$ half-filled states of both spins (provided the condition (24) is satisfied), which is entangled, with the construction of a maximum number of Bell pairs. One of the ways to remove this ambiguity in $\mathcal{H}_{\mathrm{ground}}$ is by specifying all the quantum numbers of a state. Interestingly, all the quantum numbers are furnished by the local occupation numbers, ie. $n_{i,\uparrow}n_{i,\downarrow}$, thanks to the local $U(1)$ charges. Specifying all the local site occupation numbers for each spin means fixing a particular computational basis state, which, by definition, is entanglement-free.
Everything exactly being solvable in the ultra-local model above makes it rather uninteresting. Hence, it would be curious to ask about the fate of entanglement if one turns on at least nearest neighbour hopping terms in the Hamiltonian perturbatively. The degeneracy of the $3^{L}$ dimensional 0 energy sub-space now gets lifted, as presented in the cartoon Figure 8. This is because neither the local $U(1)$ symmetry nor the supertranslation charges described above are any more there in the presence of terms like $\sum_{i,j,\sigma}t_{ij}c^{\dagger}_{i,\sigma}c_{j,\sigma}$. Only the $SU(2)\times SU(2)$ global symmetry of Hubbard model is intact. This results in the conservation of only the two quantum numbers: the total number of fermions of each spin type. Now, we are in the position of asking the question: what is the ground state of the perturbed system in the computational basis?
To answer this question, we first notice that at the lowest order in degenerate perturbation theory, each lowest band (refer to the Figure 8 right panel) energy eigenstates do belong to $\mathcal{H}_{\mathrm{ground}}$ and still has the constraint $n_{i,\uparrow}n_{i,\downarrow}=0,\forall i$ satisfied. We can expect that the ground state is still half-filled, as the half-filling condition doesn’t violate the above constraint. We performed DMRG calculations for the single chain and the coupled chain system in AB stacking to verify this. The parameters’ details are furnished in Fig. 9. We kept the hopping parameters well below $0.1U$ so that the ultra-locality is only violated perturbatively. Indeed, we checked $\sum_{i}n^{\sigma}_{i}=L/2$ for the ground state.
@@
The next level of the query is definitely the bi-partite entanglement of the ground state. However, it is subject to the constraint that the zero-hopping ground state space had: $n_{i,\uparrow}n_{i,\downarrow}=0$. This means once we choose any of the ${}^{L}C_{L/2}$ half-filled states for up-spins, the down-spins do have only one way to fill half of the lattice sites. In other words, the ground state now is a linear combination of ${}^{L}C_{L/2}$ computational basis states  [^2] This argument is valid only in the lowest order in perturbation theory of hopping parameters. However, this being strong coupling perturbation theory, the perturbative correction to the states in $\mathcal{H}_{\mathrm{ground}}$ received from excited states are doubly suppressed by a factor of $\sim\delta t/U$. . This contrasts what happens when we have a free theory, ie. $U=0$, and the ground state is a linear combination of $({}^{L}C_{L/2})^{2}$ computational basis states. Hence, a drastic drop (almost to zero) in entanglement for a large $U$ regime is expected. DMRG computation validates this argument, too, as seen in panels (b) and (d) of figure 10, respectively, for a single SSH and the coupled system in AB stacking. For the DMRG computations, we present here only the OBC case. For large $U\gg t_{1},t_{2}$, all the arguments mentioned above stand both for OBC and PBC alike. However, the vanishing of entanglement holds only if edge effects are not considered, which are substantial in the OBC. As shown in the plots 9, even in this very strong coupling regime small values of the hopping parameters still result in non-trivial edge localization in the ground state for the topologically non-trivial phase. There is a clear edge occupancy imbalance in these phases, whereas the ground state has exact half-filling in each site for the trivial phase. This non-trivial edge effect is manifest in the edge entanglement measuring $2\ln 2$ in both the single SSH and the coupled chain system, as evident in panels (a) and (c) of Fig. 10. This is what we name the remnant entropy, ie. the one remnant even in the strongly coupled ultra-local theory, perturbed by a small but topologically non-trivial hopping term.
@@ VI Conclusion and Outlook
In this work, a bi-layer SSH model with and without on-site Hubbard interaction is explored in light of the entanglement properties of the system. As mentioned earlier, for free theory, i.e. when the Hubbard interaction is absent, there is a good analytic control over the model, and one can calculate entanglement entropies from the correlation matrix of the theory semi-analytically. With a proper definition of edge entanglement, one can find a clear signature of the presence of topological order in the system, establishing the edge entanglement as an order parameter for the quantum phase transition. Within this study, we have even found that edge contribution calculated from the symmetry-resolved entanglement for the system does qualify as an order parameter. The concept of equipartition of entanglement and the study of symmetry-resolved entanglement are relatively recent concepts, and the power of these in systems with global symmetries is yet to be fully explored. Our analysis demonstrates a clear application for studying symmetry-resolved entanglement. Instead of studying the entire system, focusing only on a particular global symmetry sector is computationally cheap and would allow one to invade a class of theories which were computationally intractable otherwise. The impact of symmetry-resolved entanglement for an interacting theory and a theory containing non-Abelian anyons, would be of extreme interest to the community.
@@
Moving beyond the free theory, analytic understanding of the system is limited. Studying the spectrum and its entanglement structure for such a system is extremely expensive in terms of classical computational resources. In this work, we utilize state-of-the-art Hamiltonian simulation algorithms via the tensor network ansatz for the states and operators of the system. Note that, tensor network algorithms may not be the most efficient ones. Still, they are useful tools for studying the low energy spectrum for reasonably large systems on a classical computer, which is also essential for benchmarking quantum algorithms. Using tensor network algorithms for a fermionic system and for a system with periodic boundary conditions, are particularly nontrivial, albeit being essential for this project. Benchmarking the tensor network codes by reproducing the exact analytic results for a free theory is crucial, and has been implemented in this work. This study establishes the power of DMRG calculation that is to be used in future for more complicated systems with a rich symmetry structure. Note that, the DMRG algorithms, being governed by the study of symmetry-resolved entanglement, can access much larger system size, and do provide signatures of different topological phases and result is reliable as established in section III. Using tensor network algorithms, this work establishes the edge entanglement to be an order parameter for the topological phase transition in the case of an interacting theory.
@@
Studying an interacting theory is physically of utmost interest as it mimics the real physical systems more reliably. One important finding of this work is the notion of a remnant entanglement in the topological nontrivial phase, which remains protected for arbitrarily large value of the interaction strength. This denotes the edge entanglement to remain protected under strong correlation. This fact might impact engineering stable correlated qubits using real physical systems.
@@
In this work, the study of the remnant entropy for an interacting system led to a fascinating understanding of the basic structure of the low-energy Hilbert spaces. While, it is well understood that for arbitrary large strength of Hubbard interaction, the energy eigenstates of the ultralocal Hamiltonian are not entangled. However, even if there is a hopping present in the perturbative regime, this picture changes and the ground state acquires a quantized amount of entanglement entropy, which we refer to as the remnant entanglement and is robust under any change of the parameter values, as long as the system remains in the non-trivial topological phase. The study of the ground state configuration for interacting theories does help in revealing the mystery as discussed in detail in the last part of section VI. Future research would definitely explore this in detail for more complicated models, if possible in higher dimension. We expect the upcoming era of quantum computation will be able to shed some light on understanding the structure of the occupation number basis in real space and its sub-structures. Studying the fragmentation of the Hilbert space is important for choosing a suitable sector to work on with limited computational resources as well as for understanding underlying global symmetries for each sector.

### 08 Jan 2024 v2

@@ I Introduction
Albeit, topological insulators have been classified [10, 11] even in the presence of interaction, defining an order parameter is still not well understood. In the Su-Schrieffer-Heeger (SSH) model [12, 13, 14], the prototypical example of 1-dimensional topological insulators, two distinct topological phases are attributed due to staggered hopping amplitudes. In the non-trivial topological phase, there are two zero energy edge modes. Motivated by the fact that the existence of edge modes is a prominent signature of a topologically non-trivial phase, the entanglement entropy of edge modes has been studied in SSH or similar models extensively in the presence of interactions as well [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Following this cue, we probe a simple model with a non-trivial topological phase structure, i.e. a bi-layer SSH [26, 27, 28, 29, 30, 31] model via studying its entanglement features for a wide range of interaction strengths.
@@
We study the system for the extremely large value of interaction strength more closely in section V. If one neglects the hopping terms at very large coupling, the theory becomes ultra-local. Ultra-local models like these are called Carrollian field theories, which have recently gained interest in condensed matter [55, 56], fluid dynamics [57, 58, 59, 60], high energy [61, 62, 63] and gravitational physics [64, 65] for diverse inquiries. In particular, fermionic models with these symmetries have been dealt with in detail in [66, 67, 68]. For ultra-local systems like this, the von Neuman entropy is expected to drop to zero, which is also supported by DMRG calculations. However, if one considers including the hopping terms, albeit perturbatively small, their effects still characterize the topological phase of the system, depending upon the values of the parameters, and that is reflected in the ground state entanglement of the system as illustrated in Fig. 9.
@@ III.3 Symmetry resolved entanglement
where $\mathcal{Z}_{n}(q)$ is given by the $n$-sheets replica copy of the density matrix trace projected on the $q$-charge sector:
@@ IV.1.3 Discussion
has zero energy with respect to the free theory Hamiltonian in the non-trivial topological phase. The $A$ coefficients appearing in (16) are complex numbers subject to the condition that the state is normalized. Hence, the edge-state is 4-fold degenerate. For the same reason, the ground state,
@@
is also degenerate. Here, $b_{k}$ are all the finite, negative energy modes (holes). The 4-bilinears $\alpha^{\dagger}_{i,\uparrow}\alpha^{\dagger}_{j,\downarrow}$ are bosonic in nature. Hence, the degenerate edge state space is an orbit of $SU(4)$ action on linear combinations of 4 bosons.
At the thermodynamic limit, $\alpha_{i}$ are exactly localized at the edges, and the maximally entangled edge state is the one when all the 4 $A$’s appearing in (16) are equal, modulo $U(1)$ phases. In that case, the edge state is a couple of Bell pairs. This gives an entanglement entropy of $2\ln 2$.
@@ V Entanglement signatures at extreme coupling
Choosing a set of functions $f$ as Dirac delta functions supported at each point on the spatial line, we see that we have a conserved operator for each degree of freedom located at each spatial point, which mutually commutes. In the discretized version, these operators are $U\psi^{\dagger}_{\uparrow}\psi_{\uparrow}\psi^{\dagger}_{\downarrow}\psi_{\downarrow}$, ie the Hamiltonian density at each point. Hence, the system is exactly solvable. In fact, in this model, the local $U(1)$ phase transformations:
@@
However, the ground state being so much degenerate, the meaning of bi-partition entanglement becomes ambiguous to a large extent. For example, the vacuum state, which also is a member of the 0 energy subspace, is entanglement-free. This is a prototypical example of an entanglement-free ground state of ultra-local conformal theories [76], at least in cases where the ground state is unique. On the other hand, one can construct a linear combination of ${}^{L}C_{L/2}$ half-filled states of both spins (provided the condition (24) is satisfied), which is entangled, with the construction of a maximum number of Bell pairs. One of the ways to remove this ambiguity in $\mathcal{H}_{\mathrm{ground}}$ is by specifying all the quantum numbers of a state. Interestingly, all the quantum numbers are furnished by the local occupation numbers, i.e. $n_{i,\uparrow}n_{i,\downarrow}$, thanks to the local $U(1)$ charges. Specifying all the local site occupation numbers for each spin means fixing a particular computational basis state, which, by definition, is entanglement-free.
Everything exactly being solvable in the ultra-local model above makes it rather uninteresting. Hence, it would be curious to ask about the fate of entanglement if one turns on at least nearest neighbour hopping terms in the Hamiltonian perturbatively. The degeneracy of the $3^{L}$ dimensional 0 energy sub-space now gets lifted, as presented in the cartoon Figure 8. This is because neither the local $U(1)$ symmetry nor the supertranslation charges described above are any more there in the presence of terms like $\sum_{i,j,\sigma}t_{ij}c^{\dagger}_{i,\sigma}c_{j,\sigma}$. Only the $SU(2)\times SU(2)$ global symmetry of the Hubbard model is intact. This results in the conservation of only the two quantum numbers: the total number of fermions of each spin type. Now, we are in the position of asking the question: what is the ground state of the perturbed system in the computational basis?
To answer this question, we first notice that at the lowest order in degenerate perturbation theory, each lowest band (refer to the Figure 8 right panel) energy eigenstates do belong to $\mathcal{H}_{\mathrm{ground}}$ and still has the constraint $n_{i,\uparrow}n_{i,\downarrow}=0,\forall i$ satisfied. We can expect that the ground state is still half-filled, as the half-filling condition doesn’t violate the above constraint. To verify this, we performed DMRG calculations for the single chain and the coupled chain system in AB stacking. The parameters’ details are furnished in Fig. 9. We kept the hopping parameters well below $0.1U$ so that the ultra-locality is only violated perturbatively. Indeed, we checked $\sum_{i}n^{\sigma}_{i}=L/2$ for the ground state.
@@
The next level of the query is definitely the bi-partite entanglement of the ground state. However, it is subject to the constraint that the zero-hopping ground state space had: $n_{i,\uparrow}n_{i,\downarrow}=0$. This means once we choose any of the ${}^{L}C_{L/2}$ half-filled states for up-spins, the down-spins do have only one way to fill half of the lattice sites. In other words, the ground state now is a linear combination of ${}^{L}C_{L/2}$ computational basis states  [^2] This argument is valid only in the lowest order in perturbation theory of hopping parameters. However, this being strong coupling perturbation theory, the perturbative correction to the states in $\mathcal{H}_{\mathrm{ground}}$ received from excited states are doubly suppressed by a factor of $\sim\delta t/U$. . This contrasts what happens when we have a free theory, ie. $U=0$, and the ground state is a linear combination of $({}^{L}C_{L/2})^{2}$ computational basis states. Hence, a drastic drop (almost to zero) in entanglement for a large $U$ regime is expected. DMRG computation validates this argument, too, as seen in panels (b) and (d) of figure 10, respectively, for a single SSH and the coupled system in AB stacking. For the DMRG computations, we present here only the OBC case. For large $U\gg t_{1},t_{2}$, all the arguments mentioned above stand both for OBC and PBC alike. However, the vanishing of entanglement holds only if edge effects are not considered, which are substantial in the OBC. As shown in the plots 9, even in this very strong coupling regime small values of the hopping parameters still result in non-trivial edge localization in the ground state for the topologically non-trivial phase. There is a clear edge occupancy imbalance in these phases, whereas the ground state has exact half-filling in each site for the trivial phase. This non-trivial edge effect is manifest in the edge entanglement measuring $2\ln 2$ in both the single SSH and the coupled chain system, as evident in panels (a) and (c) of Fig. 10. This is what we name the remnant entropy, i.e. the one remnant even in the strongly coupled ultra-local theory, perturbed by a small but topologically non-trivial hopping term.
@@ VI Conclusion and Outlook
In this work, a bi-layer SSH model with and without on-site Hubbard interaction is explored in light of the entanglement properties of the system. As mentioned earlier, for free theory, i.e. when the Hubbard interaction is absent, there is good analytic control over the model, and one can calculate entanglement entropies from the correlation matrix of the theory semi-analytically. With a proper definition of edge entanglement, one can find a clear signature of the presence of topological order in the system, establishing the edge entanglement as an order parameter for the quantum phase transition. Within this study, we have even found that edge contribution calculated from the symmetry-resolved entanglement for the system does qualify as an order parameter. The concept of equipartition of entanglement and the study of symmetry-resolved entanglement are relatively recent concepts, and the power of these in systems with global symmetries is yet to be fully explored. Our analysis demonstrates a clear application for studying symmetry-resolved entanglement. Instead of studying the entire system, focusing only on a particular global symmetry sector is computationally cheap and would allow one to invade a class of theories which were computationally intractable otherwise. The impact of symmetry-resolved entanglement for an interacting theory and a theory containing non-Abelian anyons would be of extreme interest to the community.
@@
Moving beyond the free theory, analytic understanding of the system is limited. Studying the spectrum and its entanglement structure for such a system is extremely expensive in terms of classical computational resources. In this work, we utilize state-of-the-art Hamiltonian simulation algorithms via the tensor network ansatz for the states and operators of the system. Note that tensor network algorithms may not be the most efficient ones. Still, they are useful tools for studying the low energy spectrum for reasonably large systems on a classical computer, which is also essential for benchmarking quantum algorithms. Using tensor network algorithms for a fermionic system and for a system with periodic boundary conditions are particularly nontrivial, albeit essential for this project. Benchmarking the tensor network codes by reproducing the exact analytic results for a free theory is crucial and has been implemented in this work. This study establishes the power of DMRG calculation to be used in the future for more complicated systems with a rich symmetry structure. Note that the DMRG algorithms, being governed by the study of symmetry-resolved entanglement, can access much larger system size and do provide signatures of different topological phases, and the result is reliable as established in section III. Using tensor network algorithms, this work establishes the edge entanglement to be an order parameter for the topological phase transition in the case of an interacting theory.
@@
Studying an interacting theory is of utmost interest as it mimics real physical systems more reliably. One important finding of this work is the notion of a remnant entanglement in the topological nontrivial phase, which remains protected for arbitrarily large value of the interaction strength. This denotes the edge entanglement to remain protected under strong correlation. This fact might impact engineering stable correlated qubits using real physical systems.
@@
In this work, the study of the remnant entropy for an interacting system led to a fascinating understanding of the basic structure of the low-energy Hilbert spaces. It is well understood that for arbitrary large strength of Hubbard interaction, the energy eigenstates of the ultralocal Hamiltonian are not entangled. However, even if there is a hopping present in the perturbative regime, this picture changes and the ground state acquires a quantized amount of entanglement entropy, which we refer to as the remnant entanglement and is robust under any change of the parameter values, as long as the system remains in the non-trivial topological phase. The study of the ground state configuration for interacting theories does help in revealing the mystery, as discussed in detail in the last part of section VI. Future research would definitely explore this in detail for more complicated models, if possible in higher dimensions. We expect the upcoming era of quantum computation will be able to shed some light on understanding the structure of the occupation number basis in real space and its sub-structures. Studying the fragmentation of the Hilbert space is important for choosing a suitable sector to work on with limited computational resources as well as for understanding underlying global symmetries for each sector.